MaxiDrive manufactures a wide variety of parts for recreational boating, including a gear and driveshaft part for high-powered outboard boat engines. Original equipment manufacturers such as Mercury and Honda purchase the components for use in large, powerful outboards. The part sells for $694, and sales volume averages 33,500 units per year. Recently, MaxiDrive’s major competitor reduced the price of its equivalent unit to $649. The market is very competitive, and MaxiDrive realizes it must meet the new price or lose significant market share. Management has begun paying closer attention to costs and has reconfirmed the current existing standard costs. The controller then assembled the following cost and usage data for the most recent year for MaxiDrive’s production of 33,500 units:

Budgeted
Quantity
Budgeted
Cost
Actual
Quantity
Actual
Cost
Direct materials $ 7,350,000 $ 7,850,000
Direct labor 2,705,000 3,050,000
Indirect labor 2,840,000 2,665,000
Inspection (hours and cost) 1,260 470,000 1,850 435,000
Materials handling (number of purchases and cost) 6,050 925,000 4,300 910,000
Machine setups (number and cost) 2,250 1,175,000 2,350 1,150,000
Returns and rework (number of times and cost) 470 165,000 670 215,000
$ 15,630,000 $ 16,275,000

Required:

1. Calculate the target cost for maintaining current market share and profitability. (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

SOLUTION

The target cost for maintaining current market share and profitability is $440.82

Unit cost = Actual cost / Number of units

= $16,275,000 / 33,500 = 485.82

Current profit = Selling price - Unit cost

= $694 - 485.82 = 208.18

Target cost = Competitor price - Current profit

= $649 - 208.18 = $440.82